A 95.0-Ω resistor has a constant current of 3.50 A passing through it. What is the rate at which heat is being released from this resistor (in Watts)?
Answer #1

Heat released =i^2*R*t

Rate of change of heat = heat released /time

So from the above equation

Rate of change of heat = i^2*r

= 3.5*3.5*95

=1163.75 watt